# Multi-Robot Connected Fermat Spiral Coverage

### v1

@@ 7 Function PairwiseIsoverticesSplitting($\mathcal{T}_{u},u,v$):
13       $h\leftarrow$ sum of the standard deviation of the tree costs in $\mathcal{T}_{u}^{\prime}$ and the distance corresponding to any nonadjacent edge set $h^{*}$ to $h$ and $\mathcal{T}^{*}_{u}$ to $\mathcal{T}_{u}^{\prime}$ if $h<h^{*}$
@@
4.3 Optimization: MMRTC Solution Refinement
@@
4.4 Case Study: MMRTC Solution Optimizations
@@ 5 Empirical Evaluation
This section presents our experimental results on a 3.49 GHz Apple® M2 CPU laptop with 16GB RAM. Our code will be publicly available upon acceptance of this paper.
@@
Comparison: We compare MCFS (+BOTH) with two state-of-the-art grid-based MCPP methods, TMC (Vandermeulen, Groß, and Kolling 2019) and $*{}^{*}$ (Lu et al. 2023), that minimize path turns. To adapt TMC and $*{}^{*}$ to the non-rectilinear workspaces in our instances, we use overlay grids to approximate the workspaces, followed by shortest pathfinding for robot return to root positions post-coverage. Note that the reported coverage and overlapping ratios for TMC and $*{}^{*}$ are approximations due to the workspace approximation and small intersection of their coverage paths with obstacles, whereas the values for MCFS are exact. In Tab. 2, while the average coverage ratios of TMC, $*{}^{*}$, and MCFS are comparably close (with a 3.51% variance), MCFS demonstrates an average makespan reduction of 32.0% and 27.9%, curvature reduction of 75.7% and 47.8%, and overlapping ratio reduction of 13.6% and 20.9% compared to TMC and $*{}^{*}$, respectively. Both MCFS and TMC require longer runtime due to solving MIP models for MMRTC and MTSP, respectively, especially in instances with larger isographs or more robots (e.g., office). Fig. 8 and Fig. 9 visualize the coverage paths via TMC and $*{}^{*}$, respectively. These paths exhibit a back-and-forth boustrophedon pattern, leading to high curvature and imperfect coverage around complex obstacles. In contrast, MCFS notably excels in generating smooth paths that efficiently contour around arbitrarily shaped obstacles, a clear visual advantage over the other methods as shown in Fig. 6.
@@
Figure 9: $*{}^{*}$ MCPP solutions for 2-torus and office.
@@ 6 Conclusions
We proposed the MCFS framework, an innovative approach that blends principles from computer graphics and automated planning to tackle the challenges of covering arbitrarily shaped workspaces in complex MCPP tasks. Future work includes improving isoline quality to further boost the coverage ratio, incorporating kinodynamic constraints into the generation and stitching procedures of the isolines, and developing heuristics to accelerate the PIS function and the MMRTC solving for large numbers of robots or isolines.
@@ Appendix A Connected Fermat Spiral (CFS)
A.1 Identifying the Pocket Regions
@@
A.2 Generating Fermat Spiral in Pocket
@@
A.3 Connecting Separated Fermat Spirals

### v3

@@ 25        $\mathcal{T}_{u}^{\prime}\leftarrow$ a copy of $\mathcal{T}_{u}$
34
35       $h\leftarrow$ sum of the standard deviation of the tree costs in $\mathcal{T}_{u}^{\prime}$ and the distance corresponding to any nonadjacent edge
36        set $h^{*}$ to $h$ and $\mathcal{T}^{*}_{u}$ to $\mathcal{T}_{u}^{\prime}$ if $h<h^{*}$
37
@@
Optimization: MMRTC Solution Refinement
@@
Case Study: MMRTC Solution Optimizations
@@ 5 Empirical Evaluation
This section presents our experimental results on a 3.49 GHz Apple® M2 CPU laptop with 16GB RAM.
@@
Comparison: We compare MCFS (+BOTH) with two state-of-the-art grid-based MCPP methods, TMC (Vandermeulen, Groß, and Kolling 2019) and TMSTC (Lu et al. 2023), that minimize path turns. To adapt TMC and TMSTC to the non-rectilinear workspaces in our instances, we use overlay grids to approximate the workspaces, followed by shortest pathfinding for robot return to root positions post-coverage. Note that the reported coverage and overlapping ratios for TMC and TMSTC are approximations due to the workspace approximation and small intersection of their coverage paths with obstacles, whereas the values for MCFS are exact. In Tab. 2, while the average coverage ratios of TMC, TMSTC, and MCFS are comparably close (with a 3.51% variance), MCFS demonstrates an average makespan reduction of 32.0% and 27.9%, curvature reduction of 75.7% and 47.8%, and overlapping ratio reduction of 13.6% and 20.9% compared to TMC and TMSTC, respectively. Both MCFS and TMC require longer runtime due to solving MIP models for MMRTC and MTSP, respectively, especially in instances with larger isographs or more robots (e.g., office). Fig. 8 and Fig. 9 visualize the coverage paths via TMC and TMSTC, respectively. These paths exhibit a back-and-forth boustrophedon pattern, leading to high curvature and imperfect coverage around complex obstacles. In contrast, MCFS notably excels in generating smooth paths that efficiently contour around arbitrarily shaped obstacles, a clear visual advantage over the other methods as shown in Fig. 6.
@@
Figure 9: TMSTC MCPP solutions for 2-torus and office.
@@ 6 Conclusions
We proposed the MCFS framework, an innovative approach that blends principles from computer graphics and automated planning to tackle the challenges of covering arbitrarily shaped workspaces in complex MCPP tasks. Future work includes improving isoline quality to further boost the coverage ratio, as the coverage of MCFS highly depends on whether the generated isolines can adequately cover the workspace. incorporating kinodynamic constraints into the generation and stitching procedures of the isolines, and developing heuristics to accelerate the PIS function and the MMRTC solving for large numbers of robots or isolines.
@@ Appendix A Connected Fermat Spiral (CFS)
Identifying the Pocket Regions
@@
Generating Fermat Spiral in Pocket
@@
Connecting Separated Fermat Spirals